OpenStudy (anonymous):

Evaluate: f(x) = 2x^5 -3x^4 + x^3 - 6x + 20: x = 2 and g(x) = -8x^3 + 5x^2 + 14; x = -3

OpenStudy (maheshmeghwal9):

hey man put 2 in place of f(x) & by MDAS rule, try to solve it. do the same in g(x). but remember if power is even (-) becomes (+) & vice versa if odd:)

OpenStudy (maheshmeghwal9):

in place of x* in f(x)

OpenStudy (anonymous):

ok thanks

OpenStudy (maheshmeghwal9):

one I do for u\[2(2)^5-3(2)^4+2^3-6(2)+20=64-48+8-12+20= u know \it\]

OpenStudy (anonymous):

thank you :3

OpenStudy (maheshmeghwal9):

yw:)